# On structural contraction of biological interaction networks

### v3

@@ 3.2.4 Boundedness
Let ${\mathscr{N}}$ be a given a network, and let the dynamical system (4) be given with a fixed kinetics $R$. Fix a stoichiometric class $\mathscr{C}$. Then, the dynamical system (4) is said to be uniformly bounded over $\mathscr{C}$ if for each compact set $K\subset\mathscr{C}$, there exists a compact set $\tilde{K}$ such that $\varphi(t,K)\subset\tilde{K}$ for all $t\geq 0$.
@@ 3.2.5 Remark on global convergence to steady states
Recent work in contraction theory [49, 34] has also provided results on the question of convergence under the assumption of nonexpansivitiy. As has been stated originally in [49] and then in [34, Theorem 2], nonexpansivity with respect to a polyhederal norm coupled with analyticity of the vector field (which applies if all reaction rates are real-analytic functions, as is the case with mass-action and most other kinetics) implies that each trajectory is either unbounded or converges to the set of steady states. Combining the latter result with Corollary 3.13 means that the existence of a single steady state implies boundness, and hence convergence to the set of steady states. However, this result does not guarantee the uniquessness of a steady state in a stoichiometric class. For that and to get global stability, we still need robust non-degeneracy of the Jacobian, see for instance [49, Theorem 19,iii].
@@ 3.2.6 Alternative statement of nonexpansivity
$\dot{\tilde{z}}=\begin{bmatrix}S_{1}\\ D\end{bmatrix}\Gamma R(x)=\begin{bmatrix}S_{1}\Gamma R(S^{-1}\tilde{z}+{\bar{x}}))\\ 0\end{bmatrix},$
@@
Since we have $x(0)\in\mathscr{C}_{{\bar{x}}}$, then $x(t)-{{\bar{x}}}\in{\mathop{\rm Im\,}}\Gamma$ for all $t$, and we immediately get that $\tilde{z}_{2}(t)=D(x-{{\bar{x}}})=0$ for all $t$. Furthermore, let us write $S^{-1}=:[U_{1},U_{2}]$, where $U_{1}\in\mathbb{R}^{n\times(n-d)},W_{2}\in\mathbb{R}^{n\times d}$, $d=\dim({\mathop{\rm Ker\,}}(\Gamma^{T}))$.
@@
Finally we get the dynamics on the $n-d$ manifold as:
@@
In the new coordinates, the problem reduces to checking contraction of (24) with respect to the norm .$\|.\|_{BU_{1}}:\tilde{z}_{1}\mapsto\|BU_{1}\tilde{z}_{1}\|_{\infty}$. Hence, we state the following corollary.
@@
Let a network ${\mathscr{N}}$ be given. Let $S,S_{1},U_{1}$ be a matrices defined as above. Consider the ODE (24) whose Jacobian is $S_{1}\Gamma\frac{\partial R}{\partial x}U_{1}$. Consider the norm .$\|.\|_{BW_{1}}:\hat{z}_{1}\mapsto\|BU_{1}\hat{z}_{1}\|_{\infty}$ and the associated logarithmic norm $\mu_{BW_{1}}$. Then, $\mu_{BW_{1}}(S_{1}\Gamma\frac{\partial R}{\partial x}U_{1})=\mu_{B,{\mathop{\rm Im\,}}\Gamma}(\Gamma\frac{\partial R}{\partial x})\leq 0$. In other words, the logarithmic norm of the reduced system is independent of the choice of $S_{1}$ and is non-positive. In addition, consider any two solutions $\hat{\varphi}(t;\hat{\ }z_{1a}),\hat{\varphi}(t;\hat{z}_{1b})$ of (24), then the following condition holds:
$|\hat{\varphi}(t;\hat{z}_{1a})-\hat{\varphi}(t;\hat{z}_{1a})|_{BW_{1}}\leq|\hat{z}_{1a}-\hat{z}_{1b}|_{BW_{1}},~{}\mbox{for all}~{}t\geq 0.$
@@ 3.4.1 The PTM cycle (cont’d)
Let us consider again the PTM cycle (9). Using Theorem 3.4 the network is nonexpansive with respect to the norm |.$|.|_{B}$ in each stoichiometric class, where $B$ is given by (15). Figure 2 shows a numerical simulation of the distance between pairs of trajectories where it can be seen that they are approaching each other. In the next section, we will show that the system is strictly contractive over any compact set in the positive orthant.
@@ 3.4.2 An unstable system
Note, however, that the trajectories of the system can become unbounded when a steady state fails to exists. Figure 2-b) shows unbounded trajectories. Nevertheless, the distance between the trajectories never increases.
@@ 4.1 Background: siphons and persistence
We revisit the concept of a siphons, which will assist us in establishing strict contraction over compact sets. A dynamical system defined over the positive orthant is said to be persistent if trajectories that start in the interior of the positive orthant do not approach the boundary asymptotically [13]. This is also known as non-extinction since it is tantamount to asking that species are initially present in the BIN do not go extinct.
@@
Let ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$ be a network. A set $P\subset\mathscr{S}$ is said to be a siphon if a set the satisfies the following statement: $\forall X\in P$, if $X$ is a product of some reaction ${\rm\bf R}_{j}$, then $\exists X^{\prime}\in P$ such that $X^{\prime}$ is a reactant of ${\rm\bf R}_{j}$. In addition, a siphon is trivial if it contains no support of a conservation law, and it is critical otherwise.
@@
Consider a network ${\mathscr{N}}$. Consider the corresponding dynamical system (4) with given admissible kinetics $R$, and assume that it is uniformly bounded over a given a stoichiometric class $\mathscr{C}$. Assume that all siphons are trivial. Then, for any compact $K\subset\mathscr{C}\cap(0,+\infty)^{n}$, there exist a number $\varepsilon>0$ and a convex and compact $\tilde{K}$ in $[\varepsilon,+\infty)^{n}$ such that $\varphi(t,K)\in\tilde{K}$ for all $t\geq 0$.
@@
Using [11, Lemma 5], there exist $\varepsilon>0$ and $\tilde{K_{1}}$ compact in $[\varepsilon,+\infty)^{n}$ such that $\varphi(t,K)\in\tilde{K}_{1}$ for all $t\geq 0$. Our desired convex $\widetilde{K}$ is any compact and convex set that satisfies $\widetilde{K}_{1}\subset\widetilde{K}\subset[\varepsilon,\infty)^{n}$. One such set is the closed box $[\varepsilon,x_{1}^{*}]\times..\times[\varepsilon,x_{n}^{*}]$, where $x^{*}\in\widetilde{K}_{1}$ is the supremum point of $\widetilde{K}_{1}$.
@@ 4.2 Strict contraction over positive compact sets
Recall that the upper bound in (22) is $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})$ which is written in term of the partial derivatives of $R$. Therefore, we if can establish a lower bound on the partial derivatives of $R$, then we can hope to show strict contraction. However, for networks that contain bimolecular reactions, this is not generally possible since the partial derivatives of $R$ can get arbitrarily small when we get close to the boundary. Hence, our technique will be to establish contraction with respect to compact sets. If we assume the absence of trivial siphons, then Lemma 4.2 assures us that there is a uniform positive lower bound on all the states. We will show that this can be applied to few examples and then we state the general result.
@@
Unlike standard two-body binding, three-body binding [32] requires to a bridging molecule $B$ for the formation of a trimeric complex $\rm ABC$. The reaction network can be written as follows:
@@
The network admits the GLF ${\widetilde{V}}(r)=\|\Gamma r\|_{\infty}$, so $B=I$. The certificate matrices $\Lambda_{\ell},\ell=1,..,12$ can be computed via Corollary 2.15, and hence the upper bound (22) can be computed explicitly, the details are provided in §6.2. For our purposes here, (22) gives us:
@@
where the last inequality holds for all $x\in\mathbb{R}_{+}^{n}$ by assumption AK3. Therefore, this establishes that the Jacobian has a strictly negative logarithmic norm over the positive orthant. However, this falls short of proving strict contraction. Nevertheless, the situation can be remedied for compact sets by using Lemma 4.2. So, let $K\subset\mathbb{R}_{+}^{n}$ be any positive compact set. By Lemma 4.2 there exists an $\varepsilon>0$ and a convex and compact $\widetilde{K}$ such that $\varphi(t;K)\in\widetilde{K}\subset[\varepsilon,\infty)^{n}$ for all $t\geq 0$. So let $c:=\min_{x\in\widetilde{K}}\rho(x)>0$. So if $x(0),y(0)\in K\subset\mathscr{C}_{\bar{x}}$, then we get strict contraction
@@
Let us continue considering the PTM cycle. In order to bound the logarithmic norm using (22), we need to examine the expression $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}(x)\Lambda_{\ell})$. By substituting the values of $\Lambda_{\ell},\ell=1,..,6$ we get: (the dependence on $x$ is dropped for notational simplicity)
@@
Let $J$ be a square matrix. Recall that $\mu_{\infty}(J)=\max_{i}\sigma_{i}(J)$, where $\sigma_{i}(J)=[J]_{ii}+\sum_{j\neq i}|J_{ij}|$. For the matrix above, note that $\sigma_{3}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})$ and $\sigma_{5}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})$ are both identically zero. Hence, $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})=0$ for all possible choices of $\rho_{1},..,\rho_{6}$. This situation precludes using an argument similar to the previous example. Nevertheless, we can adapt a trick used in [63] by using a non-singular diagonal matrix $P$ to define a scaled norm |.$|.|_{PB}$. Hence, the equation $B\Gamma(\sum_{\ell}\rho_{\ell}Q_{\ell})=(\sum_{\ell}\rho_{\ell}\Lambda_{\ell})B\Gamma$, can be written as
@@
We are going to show that it is always possible to choose $\theta$ such that the expression above is negative over an appropriate set. So let $K$ be a positive compact set. By Lemma 4.2 there exists an $\varepsilon>0$ and a convex and compact $\widetilde{K}$ such that $\varphi(t;K)\in\widetilde{K}\subset[\varepsilon,\infty)^{n}$ for all $t\geq 0$. Define the following constant:
@@
which exists and is well-defined due to the positivity and compactness of $\widetilde{K}$. Therefore, for any choice of $\theta\in(0,\bar{\theta})$, the logarithmic norm $\mu_{\infty}\left(\sum_{\ell=1}^{s}\rho_{\ell}P\Lambda_{\ell}P^{-1}\right)$ is negative over $\widetilde{K}$, and is upper bounded by $-c$ for some $c>0$. Therefore, if $x(0),y(0)\in K\subset\mathscr{C}_{\bar{x}}$, then we get strict contraction
@@ 4.2.1 Weakly contractive networks: a structural notion
Let $J\in\mathbb{R}^{n\times n}$ be matrix that satisfies $\sigma_{i}(\Lambda)\leq 0,i=1,..,n$. Let
$S_{-}:=\{i\in{1,..,n}|\sigma_{i}(\Lambda)<0\}$
@@
$S_{0}:=\{i\in{1,..,n}|\sigma_{i}(\Lambda)=0\}\,.$
If that for each $i\in S_{0}$ there exists a series of indices $i,i_{0},..,i_{k_{i}-1}\in S_{0}$ and $i_{k_{i}}\in S_{-}$ such that $J_{ii_{0}},J_{i_{0}i_{1}},..,J_{i_{k-1}i_{k}}>0$, then $\Lambda$ is said to be weakly contractive. In addition, we call $k_{i}$ the nonexpansivity depth of index $i$. Then, let $S_{0k}\subset S_{0}$ be all the indices that have a nonexpansivity depth exactly $k$. Finally, $k:=\max_{i\in S_{0}}k_{i}$ is the maximal nonexpansivity depth of $\Lambda$. If $S_{0}=\emptyset$, then $k:=0$.
Let $\theta>0$ and a weakly contractive $\Lambda$ be given. Then we define the contractor matrix $P_{\theta}:=\prod_{\Lambda=1}^{k}P^{(\Lambda)}$, where $k$ is the maximal nonexpansivity depth of $\Lambda$. The matrices $P^{(\Lambda)}$ are diagonal matrices that are defined iteratively as follows. For $\Lambda=1$, then $p_{ii}^{(1)}=1+\theta$ if $i\in S_{-}$, and $p_{ii}^{(1)}=1$ otherwise. Next, assume we have defined the matrices $P^{(1)},..,P^{(\Lambda-1)}$. To define $P^{(j)}$, we write $p_{ii}^{(j)}=1+\theta$ if $i\in S_{-}\cup S_{01}\cup..\cup S_{0(j-1)}$, and $p_{ii}^{(j)}=1$ otherwise.
@@
for any choice if $\theta>0$. If $i\in S_{0j},1<j\leq k$, then $\sigma_{i}(P^{*}\Lambda{P^{*}}^{-1})=0$. If $i\in S_{-}$, then
@@
Note exists a small enough $\bar{\theta}>0$ such that the last inequality holds for all $i\in S_{-}$ and for all $\theta\in(0,\bar{\theta})$. Then, we define $c_{\theta}:=\max_{i\in S_{-}}\sigma_{i}(\Lambda)+\sum_{j\in S_{0}}\theta|\Lambda_{ij}|$.
For the induction step, assume that the statement holds for $\tilde{k}-1$. Then, the same argument used above can be applied to $\tilde{\Lambda}=\prod_{j=1}^{\tilde{k}-1}P^{(j)}\Lambda\left(\prod_{j=1}^{\tilde{k}-1}P^{(j)}\right)^{-1}$ where it can be immediately seen that $\tilde{S}_{-}=S_{-}\cup S_{01}\cup..\cup S_{0(\tilde{k}-1)}$ and $S_{0}=S_{0(\tilde{k})}\cup..\cup S_{0k}$.
@@
$P=\begin{bmatrix}1&0&0&0&0\\ 0&1+\theta&0&0&0\\ 0&0&(1+\theta)^{2}&0&0\\ 0&0&0&(1+\theta)^{3}&0\\ 0&0&0&0&(1+\theta)^{2}\end{bmatrix}.$
Then, Lemma 4.6 implies that $\exists\bar{\theta}>0$ such that $\forall\theta\in(0,\bar{\theta})\exists c_{\theta}>0$, where $\mu_{\infty}(PJP^{-1})<c_{\theta}<0$. For instance, $\bar{\theta}=1$ for the example above. Choosing $\theta=0.1$, for instance, gives $\mu_{\infty}(PJP^{-1})=-0.0909<0$. $\blacksquare$
@@
Let square matrices $\Lambda_{\ell},\ell=1,..,s$ be given such that $\mu_{\infty}(\Lambda_{\ell})\leq 0$, and denote $\bar{\Lambda}(\rho_{1},..,\rho_{s})=\sum_{\ell}\rho_{\ell}\Lambda_{\ell}$. Then, $\bar{\Lambda}(\rho_{1},..,\rho_{s})$ is weakly contractive for all $\rho_{1},..,\rho_{s}>0$ iff (¯$\bar{(}\rho_{1},..,\rho_{s})$ is weakly contractive for at least one positive tuple $(\rho_{1},..,\rho_{s})$.
@@
Let a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$ be given and assume that it admits a GLF ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$. We say that the network is weakly contractive if there exists matrices $\Lambda_{1},..,\Lambda_{s}$ satisfying Corollary 2.15 such that $\bar{\Lambda}=\sum_{\ell=1}^{s}\Lambda_{\ell}$ is weakly contractive.

### v4

@@ 3.2.4 Boundedness
Let $\mathscr{N}$} be a given a network, and let the dynamical system (4) be given with a fixed kinetics $R$. Fix a stoichiometric class $\mathscr{C}$. Then, the dynamical system (4) is said to be uniformly bounded over $\mathscr{C}$ if for each compact set $K\subset\mathscr{C}$, there exists a compact set $\tilde{K}$ such that $\varphi(t;K)\subset\tilde{K}$ for all $t\geq 0$.
@@ 3.2.5 Remark on global convergence to steady states
Recent work in contraction theory [49, 34] has also provided results on the question of convergence under the assumption of nonexpansivity. As has been stated originally in [49] and then in [34, Theorem 2], nonexpansivity with respect to a polyhedral norm coupled with analyticity of the vector field (which applies if all reaction rates are real-analytic functions, as is the case with mass-action and most other kinetics) implies that each trajectory is either unbounded or converges to the set of steady states.
Combining the latter result with Corollary 3.13 means that the existence of a single steady state implies boundedness, and hence convergence to the set of steady states. However, this result does not guarantee the uniqueness of a steady state in a stoichiometric class. For that and to get global stability, we still need robust non-degeneracy of the Jacobian, see for instance [49, Theorem 19,iii].
@@ 3.2.6 Alternative statement of nonexpansivity
$\dot{\tilde{z}}=\begin{bmatrix}S_{1}\\ D\end{bmatrix}\Gamma R(x)=\begin{bmatrix}S_{1}\Gamma R(S^{-1}\tilde{z}+{\bar{x}})\\ 0\end{bmatrix}$,
@@
Since we have $x(0)\in\mathscr{C}_{{\bar{x}}}$, then $x(t)-{{\bar{x}}}\in{\mathop{\rm Im\,}}\Gamma$ for all $t$, and we immediately get that $\tilde{z}_{2}(t)=D(x-{{\bar{x}}})=0$ for all $t$. Furthermore, let us write $S^{-1}=:[U_{1},U_{2}]$, where $U_{1}\in\mathbb{R}^{n\times(n-d)},{U}_{2}\in\mathbb{R}^{n\times d}$, $d=\dim({\mathop{\rm Ker\,}}(\Gamma^{T}))$.
@@
Finally we get the dynamics on the ndn-d manifold as:
@@
In the new coordinates, the problem reduces to checking contraction of (24) with respect to the norm $\|.\|_{BU_{1}}:\tilde{z}_{1}\mapsto\|BU_{1}\tilde{z}_{1}\|_{\infty}$. Hence, we state the following corollary.
@@
Let a network $\mathscr{N}$} be given. Let $S,S_{1},U_{1}$ be matrices defined as above. Consider the ODE (24) whose Jacobian is $S_{1}\Gamma\frac{\partial R}{\partial x}U_{1}$. Consider the norm $\|.\|_{B{U}_{1}}:\hat{z}_{1}\mapsto\|BU_{1}\hat{z}_{1}\|_{\infty}$ and the associated logarithmic norm $\mu_{B{U}_{1}}$. Then, $\mu_{B{U}_{1}}(S_{1}\Gamma\frac{\partial R}{\partial x}U_{1})=\mu_{B,{\mathop{\rm Im\,}}\Gamma}(\Gamma\frac{\partial R}{\partial x})\leq 0$. In other words, the logarithmic norm of the reduced system is independent of the choice of $S_{1}$ and is non-positive. In addition, consider any two solutions $\hat{\varphi}(t;{\hat{z}}_{1a}),\hat{\varphi}(t;\hat{z}_{1b})$ of (24), then the following condition holds:
$|\hat{\varphi}(t;\hat{z}_{1a})-\hat{\varphi}(t;\hat{z}_{1{b}})|_{B{U}_{1}}\leq|\hat{z}_{1a}-\hat{z}_{1b}|_{B{U}_{1}},~\mbox{for all}~t\geq 0.$
@@ 3.4.1 The PTM cycle (cont’d)
Let us consider again the PTM cycle (9). Using Theorem 3.4 the network is nonexpansive with respect to the norm $|.|_{B}$ in each stoichiometric class, where $B$ is given by (15). Figure 2 shows a numerical simulation of the distance between pairs of trajectories where it can be seen that they are approaching each other. In the next section, we will show that the system is strictly contractive over any compact set in the positive orthant.
@@ 3.4.2 An unstable system
Note, however, that the trajectories of the system can become unbounded when a steady state fails to exists. Figure 2(b) shows unbounded trajectories. Nevertheless, the distance between the trajectories never increases.
@@ 4.1 Background: siphons and persistence
We revisit the concept of siphons, which will assist us in establishing strict contraction over compact sets. A dynamical system defined over the positive orthant is said to be persistent if trajectories that start in the interior of the positive orthant do not approach the boundary asymptotically [13]. This is also known as non-extinction since it is tantamount to asking that the species that are initially present in the BIN do not go extinct.
@@
Let ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$ be a network. A set $P\subset\mathscr{S}$ is said to be a siphon if it satisfies the following statement: $\forall X\in P$, if $X$ is a product of some reaction ${\rm\bf R}_{j}$, then $\exists X^{\prime}\in P$ such that $X^{\prime}$ is a reactant of ${\rm\bf R}_{j}$. In addition, a siphon is trivial if it contains the support of a conservation law, and it is critical otherwise.
@@
Consider a network $\mathscr{N}$}. Consider the corresponding dynamical system (4) with given admissible kinetics $R$, and assume that it is uniformly bounded over a given a stoichiometric class $\mathscr{C}$. Assume that all siphons are trivial. Then, for any compact $K\subset\mathscr{C}\cap(0,+\infty)^{n}$, there exist a number $\varepsilon>0$ and a convex and compact $\tilde{K}$ in $\varepsilon,+\infty)^{n}$ such that $\varphi(t;K){\subset}\tilde{K}$ for all $t\geq 0$.
@@
Using [11, Lemma 5], there exist $\varepsilon>0$ and $\tilde{K_{1}}$ compact in $\varepsilon,+\infty)^{n}$ such that $\varphi(t;K){\subset}\tilde{K}_{1}$ for all $t\geq 0$. Our desired convex $\widetilde{K}$ is any compact and convex set that satisfies $\widetilde{K}_{1}\subset\widetilde{K}\subset[\varepsilon,\infty)^{n}$. One such set is the closed box $\varepsilon,x_{1}^{*}]\times\dots\times[\varepsilon,x_{n}^{*}]$, where $x^{*}\in\widetilde{K}_{1}$ is the supremum point of $\widetilde{K}_{1}$.
@@ 4.2 Strict contraction over positive compact sets
Recall that the upper bound in (22) is $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})$ which is written in terms of the partial derivatives of $R$. Therefore, if we can establish a lower bound on the partial derivatives of $R$, then we can hope to show strict contraction. However, for networks that contain bimolecular reactions, this is not generally possible since the partial derivatives of $R$ can get arbitrarily small when we get close to the boundary. Hence, our technique will be to establish contraction with respect to compact sets. If we assume the absence of trivial siphons, then Lemma 4.2 assures us that there is a uniform positive lower bound on all the states. We will show that this can be applied to few examples and then we state the general result.
@@
Unlike standard two-body binding, three-body binding [32] requires to a bridging molecule $B$ for the formation of a trimeric complex $\rm ABC$. The BIN can be written as follows:
@@
The network admits the GLF ${\widetilde{V}}(r)=\|\Gamma r\|_{\infty}$, so $B=I$. The certificate matrices $\Lambda_{\ell},\ell=1,\dots,12$ can be computed via Corollary 2.15, and hence the upper bound (22) can be computed explicitly, the details are provided in §6.2. For our purposes here, (22) gives us:
@@
where the last inequality holds for all $x\in\mathbb{R}_{+}^{n}$ by assumption AK3. Therefore, this establishes that the Jacobian has a strictly negative logarithmic norm over the positive orthant. However, this falls short of proving strict contraction. Nevertheless, the situation can be remedied for compact sets by using Lemma 4.2. So, let $K\subset\mathbb{R}_{+}^{n}$ be any positive compact set. By Lemma 4.2 there exists an $\varepsilon>0$ and a convex and compact $\widetilde{K}$ such that $\varphi(t;K){\subset}\widetilde{K}\subset[\varepsilon,\infty)^{n}$ for all $t\geq 0$. So let $c:=\min_{x\in\widetilde{K}}\rho(x)>0$. So if $x(0),y(0)\in K\subset\mathscr{C}_{\bar{x}}$, then we get strict contraction
@@
Let us continue considering the PTM cycle. In order to bound the logarithmic norm using (22), we need to examine the expression $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}(x)\Lambda_{\ell})$. By substituting the values of $\Lambda_{\ell},\ell=1,\dots,6$ we get: (the dependence on $x$ is dropped for notational simplicity)
@@
Let $J$ be a square matrix. Recall that $\mu_{\infty}(J)=\max_{i}\sigma_{i}(J)$, where $\sigma_{i}(J)=[J]_{ii}+\sum_{j\neq i}|{[}J_{ij}{]}|$. For the matrix above, note that $\sigma_{3}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})$ and $\sigma_{5}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})$ are both identically zero. Hence, $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})=0$ for all possible choices of $\rho_{1},\dots,\rho_{6}$. This situation precludes using an argument similar to the previous example. Nevertheless, we can adapt a trick used in [63] by using a non-singular diagonal matrix $P$ to define a scaled norm $|.|_{PB}$. Hence, the equation $B\Gamma(\sum_{\ell}\rho_{\ell}Q_{\ell})=(\sum_{\ell}\rho_{\ell}\Lambda_{\ell})B\Gamma$, can be written as
@@
We are going to show that it is always possible to choose $\theta$ such that the expression above is negative over an appropriate set. So let $K$ be a positive compact set. By Lemma 4.2 there exists an $\varepsilon>0$ and a convex and compact $\widetilde{K}$ such that $\varphi(t;K){\subset}\widetilde{K}\subset[\varepsilon,\infty)^{n}$ for all $t\geq 0$. Define the following constant:
@@
which exists and is well-defined due to the positivity and compactness of $\widetilde{K}$. Therefore, for any choice of $\theta\in(0,\bar{\theta})$, the logarithmic norm $\mu_{\infty}\left(\sum_{\ell=1}^{s}\rho_{\ell}P\Lambda_{\ell}P^{-1}\right)$ is negative over $\widetilde{K}$, and is upper bounded by c-c for some $c>0$. Therefore, if $x(0),y(0)\in K\subset\mathscr{C}_{\bar{x}}$, then we get strict contraction
@@ 4.2.1 Weakly contractive networks: a structural notion
Let ${\Lambda}\in\mathbb{R}^{n\times n}$ be a matrix that satisfies $\sigma_{i}(\Lambda)\leq 0,i=1,\dots,n$. Let
$S_{-}:=\{i\in{1,\dots,n}|\sigma_{i}(\Lambda)<0\}$
@@
$S_{0}:=\{i\in{1,\dots,n}|\sigma_{i}(\Lambda)=0\}\,.$
If for each $i\in S_{0}$ there exists a series of indices $i,i_{0},\dots,i_{k_{i}-1}\in S_{0}$ and $i_{k_{i}}\in S_{-}$ such that ${\Lambda}_{ii_{0}},{\Lambda}_{i_{0}i_{1}},\dots,{\Lambda}_{i_{k_{i}-1}i_{k_{i}}}>0$, then $\Lambda$ is said to be weakly contractive. In addition, we call $k_{i}$ the nonexpansivity depth of index $i$. Then, let $S_{0k}\subset S_{0}$ be all the indices that have a nonexpansivity depth exactly $k$. Finally, $k:=\max_{i\in S_{0}}k_{i}$ is the maximal nonexpansivity depth of $\Lambda$. If $S_{0}=\emptyset$, then $k:=0$.
Let $\theta>0$ and a weakly contractive $\Lambda$ be given. Then we define the contractor matrix $P_{\theta}:=\prod_{{j}=1}^{k}P^{(j)}$, where $k$ is the maximal nonexpansivity depth of $\Lambda$. The matrices $P^{({\lambda})}$ are diagonal matrices that are defined as follows. For j=1{j}=1, then $p_{ii}^{(1)}=1+\theta$ if $i\in S_{-}$, and $p_{ii}^{(1)}=1$ otherwise. Next, assume we have defined the matrices $P^{(1)},\dots,P^{(j-1)}$. To define $P^{(j)}$, we write $p_{ii}^{(j)}=1+\theta$ if $i\in S_{-}\cup S_{01}\cup\dots\cup S_{0(j-1)}$, and $p_{ii}^{(j)}=1$ otherwise.
@@
for any choice of $\theta>0$. If $i\in S_{0j},1<j\leq k$, then $\sigma_{i}(P^{*}\Lambda{P^{*}}^{-1})=0$. If $i\in S_{-}$, then
@@
Note exists a small enough $\bar{\theta}>0$ such that the last inequality holds for all $i\in S_{-}$ and for all $\theta\in(0,\bar{\theta})$. Then, we define $c_{\theta}:=\max_{i\in S_{-}}\left(\sigma_{i}(\Lambda)+\sum_{j\in S_{0}}\theta|\Lambda_{ij}|\right)$.
For the induction step, assume that the statement holds for $\tilde{k}-1$. Then, the same argument used above can be applied to $\tilde{\Lambda}=\prod_{j=1}^{\tilde{k}-1}P^{(j)}\Lambda\left(\prod_{j=1}^{\tilde{k}-1}P^{(j)}\right)^{-1}$ where it can be immediately seen that $\tilde{S}_{-}=S_{-}\cup S_{01}\cup\dots\cup S_{0(\tilde{k}-1)}$ and $S_{0}=S_{0(\tilde{k})}\cup\dots\cup S_{0k}$.
@@
$P^{{\theta}}=\begin{bmatrix}1&0&0&0&0\\ 0&1+\theta&0&0&0\\ 0&0&(1+\theta)^{2}&0&0\\ 0&0&0&(1+\theta)^{3}&0\\ 0&0&0&0&(1+\theta)^{2}\end{bmatrix}$.
Then, Lemma 4.6 implies that $\exists~\bar{\theta}>0$ such that $\forall\theta\in(0,\bar{\theta})\exists c_{\theta}>0$, where $\mu_{\infty}(P^{{\theta}}J(P^{{\theta}})^{-1})<c_{\theta}<0$. For instance, $\bar{\theta}=1$ for the example above. Choosing $\theta=0.1$, for instance, gives $\mu_{\infty}(P^{{\theta}}J(P^{{\theta}})^{-1})=-0.0909<0$. $\blacksquare$
@@
Let the square matrices $\Lambda_{\ell},\ell=1,\dots,s$ be given such that $\mu_{\infty}(\Lambda_{\ell})\leq 0$, and denote $\bar{\Lambda}(\rho_{1},\dots,\rho_{s})=\sum_{\ell}\rho_{\ell}\Lambda_{\ell}$. Then, $\bar{\Lambda}(\rho_{1},\dots,\rho_{s})$ is weakly contractive for all $\rho_{1},\dots,\rho_{s}>0$ iff ${\bar{\Lambda}}(\rho_{1},\dots,\rho_{s})$ is weakly contractive for at least one positive tuple $(\rho_{1},\dots,\rho_{s})$.
@@
Let a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$ be given and assume that it admits a GLF ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$. We say that the network is weakly contractive if there exist matrices $\Lambda_{1},\dots,\Lambda_{s}$ satisfying Corollary 2.15 such that $\bar{\Lambda}=\sum_{\ell=1}^{s}\Lambda_{\ell}$ is weakly contractive.